OpenStudy (aaronandyson):

Two forces have the sum of their magnitude 18N and the resultant is sqrt 228 when theyre at 120 degrees . Calculate the two forces @abhisar

OpenStudy (l):

Let \(x\) and \(18- x\) be the two forces. By vector law of addition, we know that\[\sqrt{228} = \sqrt{x^2 + (18-x)^2 + 2x(18-x)\cos(120^{\circ})}\]

OpenStudy (aaronandyson):

whats the value of cos 120? 1?

OpenStudy (l):

You should know your trigonometry for that. It is \(-1/2\).

OpenStudy (aaronandyson):

so thats x^2 + 324+x^2-36x +36x - 2x^2 *-1/2

OpenStudy (aaronandyson):

it is 16N,2N?

OpenStudy (mjdennis):

For what it's worth, you can type "cos(120 deg)" into Google. It will make you look a lot more responsible, like a good citizen of OS, if you don't have to ask for help in the basics.

OpenStudy (michele_laino):

hint: starting from the formula of @L , after a simplification, we get the subsequent quadratic equation: \[3{x^2} - 54x + 96 = 0\] please solve for \(x\)

OpenStudy (aaronandyson):

so thats 16 and 2 i posted it above

OpenStudy (michele_laino):

yes! That's right!